# SuperGrad: a differentiable simulator for superconducting processors

### v2

@@ 1 Introduction
SuperGrad can be used to optimize gate fidelities and logical error rates of quantum processors [17]. Furthermore, we foresee that our library will be useful for optimizing bosonic code processors [22, 23], where complex control schemes and scaling to more modes are ongoing challenges. Additionally, a very common task in experiments is extracting parameters from the Hamiltonian (see for example [24]). This task can also be viewed as an optimization problem, where the objective function is the difference between experimental data and simulation data. The gradients computed using SuperGrad can again make the optimization process faster and more stable. This extends previous studies [25] by the ability to fit parameters using both time dynamics data and energy spectra data. We provide an example of fitting qubit parameters from experimental data in Sec. 3.5.
@@
The remainder of this paper is structured as follows. Section II presents background information on superconducting quantum processors and introduces our approach to Hamiltonian simulation and optimization. Section III provides a hands-on tutorial for designing a quantum processor and showcases examples of SuperGrad’s capabilities. Section IV benchmarks SuperGrad performances for differentiable simulation. Section V shows the extensibility of SuperGrad and future work.
@@ 2.2 Library architecture
  • Helper processes SCGraph to generate differentiable functions for specified computations. This component includes Evolve (for evaluating time evolution) and Spectrum (for assessing static properties).
@@ 2.3 Trotterization solver
An efficient method for simulating time evolution utilizes Trotterization. This technique decomposes the time evolution operator into a sequence of local operators, facilitating the efficient computation of time evolution using 1-body and 2-body operators. Consequently, memory costs are primarily influenced by quantum states, with the memory cost scaling as $O(d^{N})$ where $N$ is the number of qudits and $d$ represents the dimension of each qudit subsystem. In an alternative perspective, the quantum state may be regarded as a high-rank state tensor. Wherein the product of a local operator and state is calculated through tensor contraction, with the time cost scaling as $O(d^{N+2})$.
@@ 2.5 Continuous adjoint method
To address the large memory overhead $O_{m}$, we explore the continuous adjoint method, a technique that significantly reduces memory costs associated with differentiating through ODE. This approach treats the backpropagation problem as the task of solving an augmented ODE for the adjoint state. We also propose the local continuous adjoint method, which in addition utilizes the locality of terms in the Hamiltonian.
@@
Figure 2: Outline of the continuous adjoint method. The black lines represent time evolution in the forward pass. After the final state $\psi_{g}$ and the objective function $\mathcal{L}$ are computed, we can use the continuous adjoint method to compute the gradient. To do that, we can evolve $\psi_{g}$ and $a_{g}$ backward in time, which means we do not need to store them during the forward pass. During the backward evolution, we can accumulate the gradient according to to Eqn. 8 and Eqn. 11.
@@ 3 Using the Library: Hands-On Examples
In this section, we provide a detailed guide on the use of SuperGrad, by using the multipath coupling scheme [38] in a fluxonium qubit chain as the example. Our initial task involves optimizing the coupling strengths in a fluxonium 3-qubit system. Subsequently, we show how to construct a longer chain of fluxonium qubits based on a frequency pattern. Finally, we optimize this frequency pattern to enhance the fidelity of the simultaneous $\mathrm{CNOT}^{\otimes 5}$ gate.
The tutorials are based on the current version of SuperGrad 0.2.1. Future releases may introduce breaking changes, and we will ensure that the documentation is kept up to date on GitHub [18].
Figure 3: Sketch of a chain-like fluxonium quantum processor with fluxonium qubits biased at the sweetspot, and multipath coupling is used between neighboring qubit pairs. (a) The fluxonium 3-qubit chain arranged in a frequency pattern is highlighted in red rectangular. Microwave pulses manipulate the grey qubit via its flux drive in Sec. 3.3. (b) 10-qubit chain-like quantum processor with simultaneous gate control. The fluxonium qubits are periodically arranged in a 3-qubit chain pattern, while their parameters have additional small deviations (see Sec. 3.4).
@@ 3.1 Create a 3-qubit chain
This section includes a hands-on example using SuperGrad. We explore a 3-qubit chain consisting of three varied fluxonium qubits (highlighted by the red-lined rectangular with label (a) in Fig. 3). Both capacitive and inductive couplings exist between qubits, configured such that the left and right pairs exhibit identical coupling strengths. The device parameters for each fluxonium are sourced from multipath coupling reference designs [38]. The time-independent part of the 3-qubit chain Hamiltonian is
@@
where $H_{f,i}$ is the Hamiltonian for individual fluxonium qubit. It takes the form as
@@
where $\hat{\varphi}_{i}$ is the phase operator, and $\hat{n}_{i}$ is the conjugate charge operator. $\varphi_{\mathrm{ext},i}$ represents the external flux, we will keep $\varphi_{\mathrm{ext},i}=\pi$ to set the fluxonium qubits at their sweetspot. The multipath coupling terms satisfy
@@
where $J_{\mathrm{C}}$ and $J_{\mathrm{L}}$ are the strength of capacitive and inductive coupling.
@@
The parameters of each qubit and their couplings are managed via Python Dictionaries and can be exported as JSON files for subsequent utilization. We now explore the structure of the dictionary used in our implementation, detailed in 1. The dictionary defines each qubit type as fluxonium and sets the external flux to bias the fluxonium at its sweetspot. In this work, we will employ the specified fluxonium parameters in 1 that form a frequency pattern. Each parameter set is uniquely labeled with distinct colors such as grey, blue and green, those colors should be declared in shared_param_mark for parameter sharing.
@@ 3.3 Workflow for Time Evolution
This subsection describes the procedure for simulating the time evolution of a quantum system using the Evolve class within the Helper package, specifically focusing on the implementation of a Cross-Resonance ($\mathrm{CR}$) pulse in a 3-qubit chain. First, we initial pulse parameters for the time evolution, as detailed in the code block 6. This step involves setting the length of the $\mathrm{CR}$ pulse and computing the effective coupling strength $J_{\mathrm{eff}}$ that is extracted from multipath coupling term Eqn. 15, as well as the detuning between the dressed $0-1$ transition frequencies of qubits 1 and 2. The driving amplitude $\epsilon_{d}$ of the pulse is then computed based on the detuning and the effective coupling strength.
@@
Subsequent to setting the initial parameters, the Evolve class is employed to simulate the time evolution of the system, detailed in the code block 7. Hyperparameters are declared for the simulation, such as truncated dimension (the number of qudit’s energy levels), parameter sharing (compute gradient with respect to the frequency pattern), and compensation options (the type of errorless single-qubit rotations before and after the time evolution operator). The unitary evolution of the $\mathrm{CR}$ pulse is computed by utilizing the eigen_basis method, which performs the time evolution simulation within the multi-qubit eigenbasis. This method employs a basis transform technique that computes the evolution within the eigenbasis while benefiting from the local Hamiltonian in the product basis, the detail is shown in Appendix D.
@@ 3.4 Device and control gradient optimization
We extend the time evolution workflow as described in Sec. 3.3, to simulate systems of ten fluxonium qubits. This is achieved by applying simultaneous control pulses to execute $\mathrm{CNOT}^{\otimes 5}$ quantum gates. SuperGrad provides a multitude of options for configuring time evolution, each affecting the simulation’s accuracy and the performance of the solver. These options include the number of time steps, and the Trotter order of the Trotterization solver, and the basis for the unitary matrix. The code template for implementing the $\mathrm{CNOT}^{\otimes 5}$ gate is presented in 9. The class MPCFluxonium is a SCGraph with many custom methods, for example, we use create_cr_pulse to add $\mathrm{CR}$ pulses in a way similar to 6. We add random deviations to the device parameters by setting add_random to True. We enable the parameter sharing so that the gradient computed in Table 2 is with respect to the parameters in the frequency pattern.
@@
Given the substantial differences in optimal learning rates for each variable [17], here we will only demonstrate a proof of principle optimization loop. We first optimize the control parameters, then perform a single-step update of the fluxonium qubit parameters $E_{C/J/L}$. Finally, we execute another control optimization with the newly updated qubit parameters and reinitialized control parameters. The multipath coupling strengths remain unchanged in this optimization. We use $1-\mathcal{F}$ as the objective function to optimize the performance of the simultaneous $\mathrm{CNOT}^{\otimes 5}$ gate. This optimization is carried out using the gradient-based algorithm L-BFGS-B. The objective function during the optimization process is shown in Fig. 5. Following the optimization of the control parameters within the original frequency pattern, the gradients with respect to the device and control parameters are detailed in Table 2.
@@ 3.5 Fitting experimental data
Figure 6: The experimental data and the fitting for the spectra measurement of a fluxonium qubit. The data are normalized as a probability distribution and then plotted as a color mesh from white to grey. The label of the x-axis is voltage, which is approximated linear with external flux $\varphi_{\text{ext}}$. After optimizing the KL-DIV objective function, we plot the result $f_{01}$ as orange dots. (a) The experiment spectra measurement with proper normalization. (b) The fitted spectra dots together with the original measured data.
@@ 4 Benchmarks
Figure 7: Benchmarking comparison of the auto-diff of Trotterization (TAD) and our local continuous adjoint method (LCAM) focused on the wall time required to evaluate (a) final state fidelity and (b) gate fidelity, as well as gradients with respect to control and device parameters, respectively. The process wall time was measured against the number of qubits, and the results were minimized across three runs. For state differentiable simulation (a), the TAD incurs a significantly higher memory cost, approximately 38 times greater than LCAM for a chain of 16 qubits (in CPU). The TAD exceeds our GPU’s memory limitation (32GB) for qubit chains larger than 13 qubits, while the LCAM method remains feasible until qubit chains exceed 24 qubits, at which point it also surpasses GPU memory limits. For unitary differentiable simulation (b), the memory cost of TAD is 30 times greater than LCAM for an 8-qubit chain (in CPU). TAD’s memory cost is out of GPU memory for qubit chains larger than 7 qubits, in contrast, the LCAM method remains within GPU memory up to a qubit chain size of 12. By performing the unitary differentiable simulation in 8GPUs configuration, the LCAM method remains available until qubit chains exceed 14 qubits. We note that this memory overhead of TAD is dependent on the time step size $\delta t$, which is chosen to be $0.01$ ns.
Our experimental setup consists of an AMD® EPYC® 7513 CPU running at 3.2 GHz (32-core/64-threads), paired with 256 GB of random-access memory (RAM), and eight NVIDIA® Tesla® V100 GPUs, each equipped with 32 GB of memory. The operating system is Ubuntu 22.04 LTS, with CUDA 12.3 and JAX 0.4.33. The SuperGrad version is 0.2.1, and the source code for the benchmarks presented in this section is accessible at [18].
@@ 4.2 Leveraging multi-GPU configuration
We have implemented multi-GPU support for time evolution unitary simulations. In our time evolution solvers, the largest memory cost is for storing quantum states array, which scales as $O(d^{2N})$. Where the quantum processor has $N$ qudits, and each qudit subsystem has a dimension $d$. By distributing the unitary across multiple devices, we can simulate larger quantum processors in parallel.

### v3

@@ 1 Introduction
SuperGrad can be used to optimize gate fidelities and logical error rates of quantum processors [17]. Furthermore, we foresee that our library will be useful for optimizing bosonic code processors [22, 23], where complex control schemes and scaling to more modes are ongoing challenges. Additionally, a very common task in experiments is extracting parameters from the Hamiltonian (see, for example [24]). This task can also be viewed as an optimization problem, where the objective function is the difference between experimental data and simulation data. The gradients computed using SuperGrad can again make the optimization process faster and more stable. This extends previous studies [25] by the ability to fit parameters using both time dynamics data and energy spectra data. We provide an example of fitting qubit parameters from experimental data in Sec. 3.5.
@@
The remainder of this paper is structured as follows.  Sec. 2 presents background information on superconducting quantum processors and introduces our approach to Hamiltonian simulation and optimization.  Sec. 3 provides a hands-on tutorial for designing a quantum processor and showcases examples of SuperGrad’s capabilities.  Sec. 4 benchmarks SuperGrad’s performance for differentiable simulation.  Sec. 5 shows the extensibility of SuperGrad and future work.
@@ 2.2 Library architecture
  • Helper processes SCGraph to generate differentiable functions for specific computations. This component includes Evolve (for evaluating time evolution) and Spectrum (for assessing static properties).
@@ 2.3 Trotterization solver
An efficient method for simulating time evolution utilizes Trotterization. This technique decomposes the time evolution operator into a sequence of local operators, facilitating the efficient computation of time evolution using 1-body and 2-body operators. Consequently, memory costs are primarily influenced by quantum states, with the memory cost scaling as $O(d^{N})$ where $N$ is the number of qudits, and $d$ represents the dimension of each qudit subsystem. From an alternative perspective, the quantum state may be regarded as a high-rank state tensor. Wherein the product of a local operator and state is calculated through tensor contraction, with the time cost scaling as $O(d^{N+2})$.
@@ 2.5 Continuous adjoint method
To address the large memory overhead $O_{m}$, we explore the continuous adjoint method, a technique that significantly reduces memory costs associated with differentiating through ODEs. This approach treats the backpropagation problem as the task of solving an augmented ODE for the adjoint state. We also propose the local continuous adjoint method, which in addition utilizes the locality of terms in the Hamiltonian.
@@
Figure 2: Outline of the continuous adjoint method. The black lines represent time evolution in the forward pass. After the final state $\psi_{g}$ and the objective function $\mathcal{L}$ are computed, we can use the continuous adjoint method to compute the gradient. To do that, we can evolve $\psi_{g}$ and $a_{g}$ backward in time, which means we do not need to store them during the forward pass. During the backward evolution, we can accumulate the gradient according to Eqn. 8 and Eqn. 11.
@@ 3 Using the Library: Hands-On Examples
In this section, we provide a detailed guide on the use of SuperGrad, by using the multipath coupling scheme [38] in a fluxonium qubit chain as the example. The two-fluxonium cross-resonance gate has been successfully realized in the laboratory [39, 40], which uses a similar architecture. Our initial task involves optimizing the coupling strengths in a fluxonium 3-qubit system. Subsequently, we show how to construct a longer chain of fluxonium qubits based on a frequency pattern. Finally, we optimize this frequency pattern to enhance the fidelity of the simultaneous $\mathrm{CNOT}^{\otimes 5}$ gate.
The tutorials are based on the current version of SuperGrad 0.2.3. Future releases may introduce breaking changes, and we will ensure that the documentation is kept up to date on GitHub [18].
Figure 3: Sketch of a chain-like fluxonium quantum processor with fluxonium qubits biased at the sweetspot, and multipath coupling is used between neighboring qubit pairs. (a) The fluxonium 3-qubit chain arranged in a frequency pattern is highlighted in the red rectangle. Microwave pulses manipulate the gray qubit via its flux drive in Sec. 3.3. (b) 10-qubit chain-like quantum processor with simultaneous gate control. The fluxonium qubits are periodically arranged in a 3-qubit chain pattern, while their parameters have additional small deviations (see Sec. 3.4).
@@ 3.1 Create a 3-qubit chain
This section includes a hands-on example using SuperGrad. We explore a 3-qubit chain consisting of three varied fluxonium qubits (highlighted by the red-lined rectangle with label (a) in Fig. 3). Both capacitive and inductive couplings exist between qubits, configured such that the left and right pairs exhibit identical coupling strengths. The device parameters for each fluxonium are sourced from multipath coupling reference designs [38]. The time-independent part of the 3-qubit chain Hamiltonian is
@@
where $H_{f,i}$ is the Hamiltonian for the individual fluxonium qubit. It takes the form as
@@
where $\hat{\varphi}_{i}$ is the phase operator, and $\hat{n}_{i}$ is the conjugate charge operator. $\varphi_{\mathrm{ext},i}$ represents the external flux. We will keep $\varphi_{\mathrm{ext},i}=\pi$ to set the fluxonium qubits at their sweetspot. The multipath coupling terms satisfy
@@
where $J_{\mathrm{C}}$ and $J_{\mathrm{L}}$ are the strengths of capacitive and inductive coupling.
@@
The parameters of each qubit and their couplings are managed via Python Dictionaries and can be exported as JSON files for subsequent utilization. We now explore the structure of the dictionary used in our implementation, detailed in 1. The dictionary defines each qubit type as fluxonium and sets the external flux to bias the fluxonium at its sweetspot. In this work, we will employ the specified fluxonium parameters in 1 that form a frequency pattern. Each parameter set is uniquely labeled with distinct colors, such as grey, blue and green. Those colors should be declared in shared_param_mark for parameter sharing.
@@ 3.3 Workflow for Time Evolution
This subsection describes the procedure for simulating the time evolution of a quantum system using the Evolve class within the Helper package, specifically focusing on the implementation of a Cross-Resonance ($\mathrm{CR}$) pulse in a 3-qubit chain. First, we initialize pulse parameters for the time evolution, as detailed in the code block 6. This step involves setting the length of the $\mathrm{CR}$ pulse and computing the effective coupling strength $J_{\mathrm{eff}}$ that is extracted from the multipath coupling term Eqn. 15, as well as the detuning between the dressed $0-1$ transition frequencies of qubits 1 and 2. The driving amplitude $\epsilon_{d}$ of the pulse is then computed based on the detuning and the effective coupling strength.
@@
Subsequent to setting the initial parameters, the Evolve class is employed to simulate the time evolution of the system, detailed in the code block 7. Hyperparameters are declared for the simulation, such as truncated dimension (the number of qudit’s energy levels), parameter sharing (compute gradient with respect to the frequency pattern), and compensation options (the type of errorless single-qubit rotations before and after the time evolution operator). The unitary evolution of the $\mathrm{CR}$ pulse is computed by utilizing the eigen_basis method, which performs the time evolution simulation within the multi-qubit eigenbasis. This method employs a basis transform technique that computes the evolution within the eigenbasis while benefiting from the local Hamiltonian in the product basis. The detail is shown in Appendix D.
@@ 3.4 Device and control gradient optimization
We extend the time evolution workflow as described in Sec. 3.3, to simulate systems of ten fluxonium qubits. This is achieved by applying simultaneous control pulses to execute $\mathrm{CNOT}^{\otimes 5}$ quantum gates. SuperGrad provides a multitude of options for configuring time evolution, each affecting the simulation’s accuracy and the performance of the solver. These options include the number of time steps, the Trotter order of the Trotterization solver, and the basis for the unitary matrix. The code template for implementing the $\mathrm{CNOT}^{\otimes 5}$ gate is presented in 9. The class MPCFluxonium is a SCGraph with many custom methods. For example, we use create_cr_pulse to add $\mathrm{CR}$ pulses in a way similar to 6. We add random deviations to the device parameters by setting add_random to True. We enable the parameter sharing so that the gradient computed in Table 2 is with respect to the parameters in the frequency pattern.
@@
Given the substantial differences in optimal learning rates for each variable [17], here we will only demonstrate a proof of principle optimization loop. We first optimize the control parameters, then perform a single-step update of the fluxonium qubit parameters $E_{C/J/L}$. Finally, we execute another control optimization with the newly updated qubit parameters and reinitialized control parameters. The multipath coupling strengths remain unchanged in this optimization. We use $1-\mathcal{F}$ as the objective function to optimize the performance of the simultaneous $\mathrm{CNOT}^{\otimes 5}$ gate. This optimization is carried out using the gradient-based algorithm L-BFGS-B [43]. The objective function during the optimization process is shown in Fig. 5. Following the optimization of the control parameters within the original frequency pattern, the gradients with respect to the device and control parameters are detailed in Table 2.
@@ 3.5 Fitting experimental data
Figure 6: The experimental data and the fitting for the spectra measurement of a fluxonium qubit. The data are normalized as a probability distribution and then plotted as a color mesh from white to gray. The label of the x-axis is voltage, which is approximately linear with external flux $\varphi_{\text{ext}}$. After optimizing the KL-DIV objective function, we plot the result $f_{01}$ as orange dots. (a) The experiment spectra measurement with proper normalization. (b) The fitted spectra dots together with the original measured data.
@@ 4 Benchmarks
Figure 7: Benchmarking comparison of the auto-diff of Trotterization (TAD) and our local continuous adjoint method (LCAM) focused on the wall time required to evaluate (a) final state fidelity and (b) gate fidelity, as well as gradients with respect to control and device parameters, respectively. The process wall time was measured against the number of qubits, and the results were minimized across three runs. For state differentiable simulation (a), the TAD incurs a significantly higher memory cost, approximately 38 times greater than LCAM for a chain of 16 qubits (in CPU). The TAD exceeds our GPU’s memory limitation (32GB) for qubit chains larger than 13 qubits, while the LCAM method remains feasible until qubit chains exceed 24 qubits, at which point it also surpasses GPU memory limits. For unitary differentiable simulation (b), the memory cost of TAD is 30 times greater than LCAM for an 8-qubit chain (in CPU). TAD’s memory cost is out of GPU memory for qubit chains larger than 7 qubits; in contrast, the LCAM method remains within GPU memory up to a qubit chain size of 12. By performing the unitary differentiable simulation in an 8GPUs configuration, the LCAM method remains available until qubit chains exceed 14 qubits. We note that this memory overhead of TAD is dependent on the time step size $\delta t$, which is chosen to be $0.01$ ns.
Our experimental setup consists of an AMD® EPYC® 7513 CPU running at 3.2 GHz (32-core/64-threads), paired with 256 GB of random-access memory (RAM), and eight NVIDIA® Tesla® V100 GPUs, each equipped with 32 GB of memory. The operating system is Ubuntu 22.04 LTS, with CUDA 12.3 and JAX 0.4.33. The SuperGrad version is 0.2.3, and the source code for the benchmarks presented in this section is accessible at [18].
@@ 4.2 Leveraging multi-GPU configuration
We have implemented multi-GPU support for time evolution unitary simulations. In our time evolution solvers, the largest memory cost is for storing the quantum states array, which scales as $O(d^{2N})$. Where the quantum processor has $N$ qudits, and each qudit subsystem has a dimension $d$. By distributing the unitary across multiple devices, we can simulate larger quantum processors in parallel.
@@
4.3 Comparison with other simulation software
In this subsection, we compare SuperGrad against similar software packages, such as SCQubits, QuTiP, and Qiskit Dynamics. We list their features in Table 3. SuperGrad stands out as the only library that integrates Hamiltonian construction and time evolution for superconducting processors. Therefore, it is capable of computing the gradients of control and device parameters in a single backpropagation. Qiskit Dynamics is the only one that supports perturbation-theory-based solvers, such as the Dyson series and Magnus expansion [47]. Qibo stands out for its advanced hardware accelerator support, which takes full advantage of GPU and multi-GPU devices [34].
To get a benchmark for the time and memory cost of these libraries, we will need to use a combination of SCQubits and QuTiP / Qiskit Dynamics to be near feature parity with SuperGrad. We again run the time evolution of $X^{\otimes 6}$ as the benchmark. The result is listed in Table 4.
SuperGrad SCQubits QuTiP Qiskit Dynamics Qibo
Version 0.2.3 4.3.0 5.1.1 0.5.1 0.2.16
Hamiltonian construction
Time evolution
Gradients of device parameters
Gradients of control parameters Pre-Alpha
GPU acceleration Pre-Alpha
Multi-GPU support
Table 3: A comparison of the capabilities of SuperGrad and other packages reveals notable differences. All tools were benchmarked using the version listing in the table. SuperGrad stands out by providing the full simulation pipeline, from superconducting processor Hamiltonian construction to time evolution, whereas SCQubits is limited to Hamiltonian construction. To use the time evolution solver of QuTiP and Qiskit Dynamics, one needs to supply the Hamiltonians, e.g., from SCQubits. This entire process simulation support in SuperGrad enables efficient differentiable simulation, while Qiskit Dynamics and Qibo only support backpropagation for the time evolution part. QuTiP provides partial support for differentiable simulations and GPU acceleration via optional plug-ins (QuTiP-JAX and QuTiP-CuPy), but these features are in the pre-alpha stage. Both Qibo and SuperGrad support distributed time evolution across multiple GPUs, which is crucial for simulating larger quantum systems. Additionally, Qiskit Dynamics offers perturbation-theory-based solvers, which deliver a significant speed advantage over traditional solvers, though at the cost of increased initial compilation time [47].
SuperGrad SCQubits + Qiskit Dynamics SCQubits + QuTiP
ODE solver Trotterization Dormand-Prince
Differential method LCAM CAM FDM
Gradient computation scope Device and control Control Device and control
Time Forward $f$ 2.02 s 9.28 s 9.28 s 173.81 s
Gradients $\nabla f$ 8.55 s 14.73 s 554.66 s 10151.47 s
Overhead $O_{t}$ 4.23 1.59 59.77 58.41
Memory Forward $f$ 2.01 GB 1.96 GB 1.96 GB 1.83 GB
Gradients $\nabla f$ 2.74 GB 2.97 GB 8.54 GB 6.31 GB
Overhead $O_{m}$ 1.36 1.52 4.36 3.45
Table 4: Benchmark of the performance of SuperGrad and other packages, focusing on the wall time and memory usage required to compute the simultaneous $\mathrm{X}^{\otimes 6}$ gate fidelity and its gradient. For the case that the gradient computation scope is "control", gradients are computed only for control parameters (24 parameters in total). In contrast, in the cases where the gradient computation scope is "device and control", gradients are computed for both control and device parameters (58 parameters in total). Time and memory overheads are calculated relative to the corresponding forward simulation. The Dormand-Prince solver in Qiskit Dynamics supports the continuous adjoint method, which allows for efficient backpropagation of time evolution with a small time overhead, as it bypasses gradient computation for the device Hamiltonian. The case is labeled as FDM using the standard finite difference method (FDM), where time overheads are approximated to the number of parameters. Due to Python’s garbage collection strategy, the memory usage is likely overestimated.
For fairness, the ODE solver and its error control parameters (atol = $1\times 10^{-8}$, rtol = $1\times 10^{-8}$) were kept the same for both QuTiP and Qiskit Dynamics. We used the JAX backend of Qiskit Dynamics to perform computations. For SuperGrad, we employ a time step size of $\delta t=0.01$ ns and utilize second-order Trotter decomposition. We observed the unitary infidelity $1-\mathcal{F}$ between SuperGrad and Qiskit Dynamics is smaller than $1\times 10^{-6}$. Note that this does not imply that the results they computed have the same accuracies. Therefore, the time overheads for computing gradients are more important than the absolute times in Table 4.
All benchmarks were run on the CPU since some of these packages do not have GPU backends. The minimal wall time across three runs was recorded, and memory usage was defined by the maximum memory resident set size, as reported by the psutil library during computation. Due to Python’s lazy garbage collection strategy, unused memory may not be released immediately, which affects the accuracy of memory usage benchmarks. This can lead to an overestimation of memory consumption, particularly in FDM computations. In theory, the memory usage for FDM computation should be equivalent to that of the corresponding forward computation.
One of SuperGrad’s key contributions is its support for the differentiable simulation of both control and device parameters. On the other hand, Qiskit Dynamics can only compute gradients with respect to control parameters. We can still estimate the gradients with respect to the device parameters using the finite difference method (FDM). For FDM, we used forward differences, which require one additional computation per parameter, with a step size of $1\times 10^{-4}$ for finite differences. The results of the computation wall time and memory usage are shown in Table 4. These benchmarks demonstrate that SuperGrad computes gradients more efficiently than similar software, particularly when dealing with a large number of parameters.